# Listen, Disentangle, and Control: Controllable Speech-Driven Talking Head Generation

### v1

@@ 4.1. Dataset
We use Voxceleb (Nagrani et al., 2017), MEAD (Wang et al., 2020), CREMA-D (Cao et al., 2014) and HDTF (Zhang et al., 2021a) datasets on our experiments. Voxceleb contains real world videos with large variation in face pose and occlusion of faces and audio. MEAD is a high-quality emotional audiovisual dataset that includes 60 actors/actresses and eight emotion categories. CREMA-D is a dataset with over 7,000 emotion category audios. HDTF is a high-resolution in-the-wild audio-visual dataset. Following (Siarohin et al., 2019), we crop faces from the videos and resize them to $256\times 256$. Faces move freely within a fixed bounding box and no need to align. The videos are sampled at the rate of 30 FPS and the audios are pre-processed to 16kHz. We train our IRFD on the Voxceleb and MEAD datasets with pose, emotion and identity. We train our SPEAK model on the Voxceleb and MEAD datasets. For evaluation, the test set contains videos from the MEDA, CREMA-D and HDTF dataset, which are invisible during training.
4.2. Implementation Details
We employ Adam optimizer for training. The $E_{i}$, $E_{e}$ and $E_{p}$ structure is a ResNet50 (He et al., 2016) are trained for 6 hours on eight-16GB NVIDIA Tesla V100 GPUs with a learning rate of 0.0001. Before training the SPEAK, the IRFD is trained on the Voxceleb dataset and MEAD dataset with $\mathcal{L}_{disentanglement}$. Then the generator $G_{d}$ in IRFD is discarded, while three encoders are used in the global generator $G_{g}$ training process. Our work is trained for 20 hours on eight-16GB NVIDIA Tesla V100 GPUs with a learning rate of 0.0001.
4.3. Evaluation
4.3.1. Evaluation Metrics.
@@ 4.3. Evaluation
We use peak signal-to-noise ratio (PSNR) and structural similarity (SSIM) (Wang et al., 2004) to compare the image quality of the generated video frames to the original video frames. To account for the accuracy of mouth forms and lip sync, we employ both the landmark distance (LMD) surrounding the mouths (M-LMD) (Chen et al., 2019) and the confidence score (Syncconf(Chung and Zisserman, 2017b) to measure audiovisual synchronization. Then we use the LDM on the whole face (F-LMD) to measure the accuracy of facial emotions and pose.
4.3.2. Comparing Methods.
We compared our SPEAK to those of other SOTA methods: Wav2Lip (Prajwal et al., 2020), ATVG (Chen et al., 2019), PC-AVS (Zhou et al., 2021) and EAMM (Ji et al., 2022). In our experiments, we evaluate our method in a self-driven setting on a separate test set that includes identities, emotions and poses that were not encountered during training. Notably, due to its focus on mouth area movements, Wav2Lip samples had fixed head poses. Conversely, we derived poses from videos when evaluating other methods. This methodology ensured consistency in our evaluations and allowed us to make meaningful comparisons between the different approaches. All methods are compared under the five metrics mentioned above. For M-LMD/F-LMD the lower the better, and the higher the better for other metrics.
Table 2. User studies on the MEAD dataset  (Wang et al., 2020) and CREMA-D dataset  (Cao et al., 2014) are measured using mean scores.
Method/Score MEAD CREMA-D
Degree of Lip-Sync $\uparrow$ Head Naturalness $\uparrow$ Video Realness $\uparrow$ Degree of Lip-Sync $\uparrow$ Head Naturalness $\uparrow$ Video Realness $\uparrow$
ATVG 3.89 3.71 3.82 4.23 3.54 3.85
EAMM 3.92 1.63 1.86 3.97 1.56 1.86
PC-AVS 3.41 1.87 2.33 3.42 1.86 2.45
Wav2Lip 4.03 1.42 1.54 4.21 1.31 1.58
Ground Truth 4.76 4.91 4.89 4.88 4.89 4.96
Ours 4.12 3.97 4.09 4.23 3.67 4.11
4.4. Experiment Results
4.4.1. Quantitative Comparisons.
@@ 4.4. Experiment Results
The results of the quantitative evaluation are reported in Table 1. Our method achieves the best performance among most metrics on MEAD and HDTF datasets. Since Wav2Lip merely generated lip-sync and does not change other parts of the reference images, it obtains the highest Syncconf on MEAD. The obtained score is even higher than that of the ground truth (original video frame). This can be attributed to the fact that Wav2Lip is trained with SyncNet as a discriminator, making it reasonable for Wav2Lip to achieve the highest confidence score of SyncNet on MEAD. Our confidence score of SyncNet is closest to ground truth on MEAD and the highest on HDTF dataset. Besides, our method achieves the best performance under the F-LMD and M-LMD on HDTF dataset. These findings indicate the superior quality of our generated real face (as measured by PSNR and SSIM) as well as a highly accurate lip synthesis (as measured by M-LMD, Syncconf) and facial emotions and pose (as measured by F-LMD).
4.4.2. Qualitative Comparisons.
@@
As shown in Figure 4, we make comparisions with other methods on various frames. The Identity, Reference Source (video frames after the fusion of emotion and pose) and Audio are all invisible during training. It is evident that our SPEAK can achieve high-fidelity emotions and posture styles matching the Reference Source, as well as more accurate lip shapes. In terms of lip-sync, Wav2lip and ATVG are competitive with our method. However, Wav2lip and ATVG do not consider emotions, so they generate seemingly reasonable lip movements but only with neutral emotions. EAMM and PC-AVS cannot achieve accurate lip-sync. In addition, the PC-AVS is not robust enough to data with background variations, leading the generated backgrounds are very different from the female identity (right side of Figure 4). Only EAMM and our method can achieve emotion and pose control. But EAMM can only control the emotions of the upper part of the face, while failing to control the mouth shape of expressions. Furthermore, EAMM connot preserve the speaking identity well. In contrast, our method can imitate emotional talking head from reference source driving video clips while achieving a accurate lip-sync, satisfactory identity preservation, and producing stable backgrounds.
Table 3. Ablation study on quantitative comparison of IRFD on the MEAD dataset  (Wang et al., 2020) and HDTF dataset  (Zhang et al., 2021a).
Method / Score MEAD HDTF
@@
w/o $\mathcal{L}_{identity}$ 0.63 25.23 4.26 3.29 3.06 0.66 25.47 4.11 3.13 3.01
w/o $\mathcal{L}_{cls}$ 0.75 28.11 2.78 3.27 3.81 0.79 28.33 2.65 3.08 3.82
w/o $\mathcal{L}_{pose}$ 0.65 25.89 3.39 3.51 3.72 0.67 26.12 3.26 3.27 3.68
w/o $\mathcal{L}_{emotion}$ 0.69 26.74 3.21 3.14 3.29 0.71 26.84 3.12 3.02 3.29
w/o $\mathcal{L}_{self}$ 0.58 25.45 3.87 3.66 3.12 0.61 25.79 3.79 3.54 3.07
Ours (Full Model) 0.81 29.31 2.48 3.25 3.99 0.83 29.46 2.32 3.02 3.96
Table 4. Ablation study on quantitative comparison of SPEAK on the MEAD dataset  (Wang et al., 2020) and HDTF dataset  (Zhang et al., 2021a).
Method / Score MEAD HDTF
[…]
w/o IRFD 0.55 24.77 5.25 3.88 2.96 0.58 24.89 5.17 3.71 2.92
w/o $\mathcal{L}_{GAN}$ 0.61 26.69 4.12 3.81 3.14 0.62 26.88 3.99 3.67 3.02
w/o $\mathcal{L}_{vgg}$ 0.72 27.01 3.37 3.76 3.67 0.75 27.13 3.21 3.64 3.61
w/o $\mathcal{L}_{sync}$ 0.72 28.41 2.64 4.14 3.16 0.76 28.57 2.52 3.98 3.15
Ours (Full Model) 0.81 29.31 2.48 3.25 3.99 0.83 29.46 2.32 3.02 3.96
Figure 5. Visualized results of IRFD effectiveness.
4.4.3. User Studies.
We conduct a user studies with 20 participants to gather their feedback on 30 videos generated by SPEAK and the other SOTA methods based on the MEAD dataset and the CREMA-D dataset. Users are asked to score each video on a scale of 1 to 5 based on the (1) degree of lip-sync, (2)naturalness of head motions, and (3) video realness. A score of 1-2 indicates that the generated video has a significant difference from the mouth GT and is of poor quality. A score of 3 indicates that the quality is mediocre, neither good nor bad. A score of 4-5 indicates that the generated video has good lip-sync, natural head motions, and realness. The higher the score, the better the performance. We processed the collected feedback scores by taking the mean value. As shown in Table 2, users give the highest marks on each aspect of our method on both two datasets. ATVG is highly competitive in terms of degree of Lip-Sync and naturalness of head motions, but its video realness is not perfect and is not as robust as our method on different datasets. While the naturalness of head motions in the other three methods is quite poor.
4.4.4. The Effectiveness of Disentanglement Module
We firstly use pre-trained IRFD to reconstruct real facial images by Gd without swapping operations as original comparison objects.Then we independently swap the latent codes of three facial features and reconstruct them into fake facial images. As visualized results shown in Figure 5, it can be intuitively seen that our proposed IRFD is effective in decoupling facial features, which is similar to the mapping network in styleGAN. And more quantitative comparisons results are discussed in the ablation study of Sec.4.4.5.
4.4.5. Ablation study.
We compare the emotional latent space obtained with and without IRFD training of the network on the MEAD dataset and the HDTF dataset to evaluate the performance of the disentanglement module. Additionally, we conduct ablation experiments on the overall SPEAK model. Table 3 demonstrates that overall, the full model exhibits significant advantages across all performance metrics, whether on the MEAD dataset or the HDTF dataset. When $\mathcal{L}_{identity}$ is removed, F-LMD and Syncconf metrics show the worst performance, indicating that $\mathcal{L}_{identity}$ effectively aligns facial features and plays a crucial role in synchronizing lip movements with audio signals. $\mathcal{L}_{cls}$, $\mathcal{L}_{pose}$, and $\mathcal{L}_{emotion}$ have certain impacts on facial emotion and movements. However, the removal of $\mathcal{L}_{self}$ significantly affects the similarity of generated images. Table 4 shows that the IRFD module contributes to the decoupling of information from images. When the IRFD module is removed from the full model, the quantitative findings are less accurate. This implies that the IRFD module contribute to the SPEAK. We also calculated the quantitative outcomes to assess the effectiveness of the losses introduced Sec.3.3. Each loss, as shown in Table 4, contributes to the component. When $\mathcal{L}_{sync}$ is removed, the results in M-LMD/F-LMD, SSIM, and PSNR are worse than the our method, indicating the necessity of $\mathcal{L}_{sync}$ in generating more accurate lip movements. We present the visual improvement of the refinement in Figure 6. The performance of our method exhibits satisfactory identity preservation, rich emotions, precise lip-sync, and posture. While the performance after removing $\mathcal{L}_{vgg}$ is also quite competitive, our method imitates the eye angle in the pose source more accurately.
Figure 6. Visualized results of SPEAK ablation study.
5. Conclusion
@@
In this paper, we propose a technique for generating accurately lip-synched emotional talking head with free pose and emotion control from other videos. We design a novel disentanglement module IRFD for decomposing the input sample into emotion, identity, and pose embedding. Then, to generate the talking head, we offer a novel talking head generation framework SPEAK. Qualitative and quantitative experiments indicate that our method performs very robustly in challenging scenarios, such as significant pose and emotional expression variations.

### v3

@@ III Experiment
Datasets. We use Voxceleb [23], MEAD [20], CREMA-D [24] and HDTF [21] datasets on our experiments. Voxceleb contains real world videos with large variation in face pose and occlusion of faces and audio. MEAD is a high-quality emotional audiovisual dataset that includes 60 actors/actresses and eight emotion categories. CREMA-D is a dataset with over 7,000 emotion category audios. HDTF is a high-resolution in-the-wild audio-visual dataset. We crop faces from the videos and resize them to $256\times 256$. We train our IRFD on the Voxceleb and MEAD datasets and train other SPEAK modules on the Voxceleb and MEAD datasets.
Implementation Details. We employ Adam optimizer for training. The structure of encoders in IRFD is ResNet50 [25]. Before training the SPEAK, the encoders and generator of IRFD are trained to encode facial features of pose, emotion, and identity into latent spaces. We randomly swap one type of facial feature code and fed into the IRFD generator to reconstruct full face. Then IRFD generator is discarded, while its encoders are used in the global generator $G_{g}$ training process. All models are trained on 8 NVIDIA V100 GPUs.
@@
Evaluation Metrics and Baselines. We use peak signal-to-noise ratio (PSNR) and structural similarity (SSIM) [26] to compare the image quality of the generated video frames to the original video frames. To account for the accuracy of mouth forms and lip sync, we employ both the landmark distance (LMD) surrounding the mouths (M-LMD) [3] and the confidence score (Syncconf[16] to measure audiovisual synchronization. Then we use the LDM on the whole face (F-LMD) to measure the accuracy of facial emotions and pose.We compared our SPEAK to ATVG [3], EAMM [12], SOTA GAN-based method TH-PAD [22] and Wav2Lip [4].
@@
Quantitative Evaluation. As results shown in Table I, our method outperforms most metrics across both datasets. Since Wav2Lip [4] merely generated lip-sync and does not change other parts of the reference images, it obtains the highest Syncconf on MEAD. And our Syncconf is closest to ground truth on MEAD and the highest on HDTF dataset. Besides, our method significantly outperforms the SOTA GAN-based method TH-PAD [22]. These indicates the superior quality of our generated real face (measured by PSNR and SSIM) as well as a highly accurate lip synthesis (measured by M-LMD and Syncconf) and emotions and pose (measured by F-LMD).
@@
Qualitative Evaluation. Result in Fig. 2 is evident that our SPEAK can achieve high-fidelity emotions and posture styles matching the Reference Source, as well as more accurate lip shapes. In terms of lip-sync, Wav2Lip [4] and ATVG [3] don’t consider emotions, so they generate seemingly reasonable lip movements but only with neutral emotions. EAMM [12] and TH-PAD [22] cannot achieve accurate lip-sync and pose. In contrast, SPEAK can imitate emotional talking head from reference source driving video clips while achieving a accurate lip-sync, satisfactory identity preservation.
TABLE II: User study on CREMA-D [24] are measured using mean scores.
Method / Score CREMA-D
Degree of Lip-Sync $\uparrow$ Head Naturalness $\uparrow$ Video Realness $\uparrow$
ATVG [3] 4.21 3.51 2.58
EAMM [12] 3.97 1.56 1.86
TH-PAD [22] 3.71 3.56 3.85
Wav2Lip [4] 3.82 1.31 1.58
Ours 4.23 3.67 4.11
Ground Truth 4.88 4.89 4.96
User Study. We conduct an user study with 20 participants to gather human feedback on 30 videos generated by SPEAK and other methods. After shuffling generated videos, users need to score each video on a scale of 1 (bad) to 5 (good) based on (1) degree of lip-sync, (2)naturalness of head motions, and (3) video realness. Users give the highest marks on each aspect of our method in Table II. TH-PAD [22] is highly competitive in terms of video realness and naturalness of head motions, but its lip-sync performance is worst. The head naturalness and video realness of both Wav2Lip [4] and EAMM [12] are relatively poor, while the video realness of ATVG [3] is not perfect compare to our proposed method.
Ablation study. We provide a detailed analysis of the impact of each module on the MEAD dataset. Quantitative results for each module of IRFD in Table III indicate that the IRFD Identity encoder contributes the most to SSIM and PSNR, the IRFD Pose encoder contributes most to F-LMD, and the IRFD Emotion encoder contributes most to M-LMD and Syncconf. This implies that IRFD module achieves our design motivation. The visualized results in Fig. 3 show the impact of each module in the entire SPEAK. When $\mathcal{L}_{sync}$ is removed, it is difficult to generate faces synchronized with mouth GT. Although the performance after removing $\mathcal{L}_{vgg}$ or $\mathcal{L}_{GAN}$ is also quite competitive, our method imitates the eye angle in the pose source more accurately. When IRFD module is removed, the less accurate qualitative findings are hard to achieve better pose and emotion control.
TABLE III: Ablation study on quantitative comparison on MEAD [20].
Method / Score MEAD
@@
w/o $\text{IRFD}_{\text{Identity}}$ 0.53 23.89 4.22 3.84 4.11
w/o $\text{IRFD}_{\text{Pose}}$ 0.57 24.77 5.15 3.78 3.98
w/o $\text{IRFD}_{\text{Emotion}}$ 0.62 25.31 3.17 4.72 3.23
Ours (Full Model) 0.84 29.41 2.48 2.92 4.99
Figure 3: Visualized results of SPEAK ablation study.
IV Conclusion
@@
In this paper, we propose a technique to generate accurately lip-synched emotional talking head with free pose and emotion control from other videos. We design a novel disentanglement module IRFD for decomposing input sample into emotion, identity, and pose embedding. Then we provide a novel talking head generation framework SPEAK. Qualitative and quantitative experiments indicate that our method performs very robustly in challenging scenarios, such as significant pose and emotional expression variations.